Subscribe: Zen Internet Service Alerts
http://status.zensupport.co.uk/rss/active.rss2.xml
Added By: Feedage Forager Feedage Grade A rated
Language: English
Tags:
aug edited  aug status  aug  edited  jul  maintenance  regret inconvenience  start  status  tue aug  tue jul  tue  wed aug  wed  zen 
Rate this Feed
Rate this feedRate this feedRate this feedRate this feedRate this feed
Rate this feed 1 starRate this feed 2 starRate this feed 3 starRate this feed 4 starRate this feed 5 star

Comments (0)

Feed Details and Statistics Feed Statistics
Preview: Zen Internet Service Alerts

Zen Internet Service Alerts



This RSS feed provides an overview of the current status of all active service faults reported by Zen Internet.



Published: Wed, 16 Aug 2017 21:37:34 +0100

Last Build Date: Wed, 16 Aug 2017 21:37:34 +0100

 



#4922 Shared Windows Hosting - Plesk emergency maintenance - mssql01a.plesk (New)

Tue, 15 Aug 2017 08:50:26 +0100

Our engineers are currently carrying out emergency maintenance, mssql01a.plesk service will be unavailable for 5-10 minutes to perform emergency maintenance. Zen regret any inconvenience this may cause. Start: Tue, 15th Aug 2017 08:46 End: Tue, 15th Aug 2017 09:00 Edited: Tue, 15th Aug 2017 08:50 Status: Outage Maintenance: Emergency



#4914 Zen Web Sites - Planned Maintenance - Secure File Sending (Close)

Mon, 14 Aug 2017 15:04:36 +0100

Maintenance completed Start: Sun, 13th Aug 2017 21:00 End: Sun, 13th Aug 2017 23:55 Clear: Mon, 14th Aug 2017 00:00 Edited: Mon, 14th Aug 2017 15:04 Status: Up Maintenance: None



#4921 Zen Web Sites - Web Site Maintenance (New)

Mon, 14 Aug 2017 13:28:08 +0100

Due to planned maintenance the following areas of the Zen website will be unavailable between Sunday Sep 10th, 2017 8.30pm -11pm : Customer Portal (https://portal.zen.co.uk & https://myaccount.zen.co.uk) Partner Portal (https://partner.zen.co.uk & https://partner2.zen.co.uk) https://secure.zen.co.uk Data Centre Access Management Zen regret any inconvenience this may cause. Start: Sun, 10th Sep 2017 20:30 End: Sun, 10th Sep 2017 23:00 Edited: Mon, 14th Aug 2017 13:28 Status: Outage Maintenance: Planned



#4920 Zen Web Sites - Partner Portal Maintenance. (New)

Mon, 14 Aug 2017 13:22:46 +0100

On the 15th August between 20:00 and 00:30 we will be conducting systems maintenance on our portal. During this maintenance window it will not be possible to modify the passwords of live Broadband services via the Modify Connection Details function in the Service Details page. After the maintenance window the function should perform as normal. Zen regret any inconvenience this may cause. Start: Tue, 15th Aug 2017 20:00 End: Wed, 16th Aug 2017 00:30 Edited: Mon, 14th Aug 2017 13:22 Status: Partial Maintenance: Planned



#4919 Shared Linux Hosting - shcp19.hosting.zen.net.uk - Maintenance (New)

Mon, 14 Aug 2017 12:47:48 +0100

Our engineers will be performing essential maintenance on the shcp19.hosting.zen.net.uk server. A short outage period is expected during this time as services are brought up to date and restarted. Zen regret any inconvenience this may cause. Start: Wed, 16th Aug 2017 22:00 End: Thu, 17th Aug 2017 00:59 Update: Thu, 17th Aug 2017 00:59 Edited: Wed, 16th Aug 2017 21:37 Status: Partial Maintenance: Planned



#4916 Zen Web Sites - www.zen.co.uk Down (Close)

Tue, 08 Aug 2017 18:48:38 +0100

Further monitoring has not found any ongoing issues. Closing ticket Start: Wed, 2nd Aug 2017 13:56 Clear: Tue, 8th Aug 2017 18:48 Edited: Tue, 8th Aug 2017 18:48 Status: Up Maintenance: None



#4917 E-mail - SPF Validation (Close)

Tue, 08 Aug 2017 09:58:34 +0100

Confirmed functioning service. Incident Closed. Start: Thu, 3rd Aug 2017 15:00 End: Tue, 8th Aug 2017 10:00 Clear: Tue, 8th Aug 2017 09:58 Edited: Tue, 8th Aug 2017 09:58 Status: Up Maintenance: None



#4918 Shared Windows Hosting - Plesk server maintenance (Close)

Tue, 08 Aug 2017 05:40:59 +0100

Maintenance successfully completed. Start: Mon, 7th Aug 2017 23:00 End: Tue, 8th Aug 2017 00:00 Clear: Wed, 9th Aug 2017 00:00 Edited: Tue, 8th Aug 2017 05:40 Status: Up Maintenance: None



#4918 Shared Windows Hosting - Plesk server maintenance (New)

Mon, 07 Aug 2017 16:04:22 +0100

Our engineers are carrying out some routine maintenance on this server, which will require a reboot. Services will be offline during this time, but should return shortly afterwards. Start: Mon, 7th Aug 2017 22:00 End: Mon, 7th Aug 2017 23:00 Edited: Mon, 7th Aug 2017 16:13 Status: Outage Maintenance: Planned



#4917 E-mail - SPF Validation (New)

Thu, 03 Aug 2017 15:49:13 +0100

From Thursday 3 August Zen's mail servers will be honoring SPF records as published by sending domains. Where SPF experiences a SPF FAIL we will reject this email. If emails are getting rejected for failing SPF we advise to check and confirm your SPF records are adequate for your domain. For domains held with Zen please call our Support teams who can offer assistance. Start: Thu, 3rd Aug 2017 15:00 End: Tue, 8th Aug 2017 10:00 Update: Tue, 8th Aug 2017 10:00 Edited: Thu, 3rd Aug 2017 15:49 Status: Up Maintenance: None



#4916 Zen Web Sites - www.zen.co.uk Down (Update)

Wed, 02 Aug 2017 17:27:36 +0100

Services to www.zen.co.uk have now been restored. All systems are currently working but we are also aware of intermittent problems. This will remain the case until our engineers have finished their investigations and have fully resolved. Further updates will be posted here when available. Zen regret any inconvenience this may cause. Start: Wed, 2nd Aug 2017 13:56 Update: Thu, 3rd Aug 2017 09:00 Edited: Wed, 2nd Aug 2017 17:27 Status: Partial Maintenance: None



#4916 Zen Web Sites - www.zen.co.uk Down (Update)

Wed, 02 Aug 2017 15:29:50 +0100

We are still working on resolving this issue. Zen regret any inconvenience this may cause. Start: Wed, 2nd Aug 2017 13:56 Update: Wed, 2nd Aug 2017 17:30 Edited: Wed, 2nd Aug 2017 15:29 Status: Outage Maintenance: None



#4916 Zen Web Sites - www.zen.co.uk Down (New)

Wed, 02 Aug 2017 14:28:43 +0100

We are currently aware of an outage with our website, www.zen.co.uk. Our engineers are investigating and further updates will be posted here when available. Please be aware that there will be no access to any customer portals at this moment. Zen regret any inconvenience this may cause. Further updates will be posted here when available. Start: Wed, 2nd Aug 2017 13:56 Update: Wed, 2nd Aug 2017 15:30 Edited: Wed, 2nd Aug 2017 14:28 Status: Outage Maintenance: None



#4915 Broadband (xDSL) - DSL Usage Records (Close)

Wed, 26 Jul 2017 22:11:09 +0100

The issue has been resolved, However customers may notice a sharp rise in usage data, as the servers catch up with the backlog. Zen regret any inconvenience this may cause. Start: Tue, 25th Jul 2017 15:00 Clear: Wed, 26th Jul 2017 22:09 Edited: Wed, 26th Jul 2017 22:11 Status: Up Maintenance: None



#4915 Broadband (xDSL) - DSL Usage Records (New)

Wed, 26 Jul 2017 13:31:06 +0100

Please be aware due to a problem on our accounting server the usage records have become backlogged. You may see on your usage graphs all of your usage over the past 24-36 hours come through at once. Zen regret any inconvenience this may cause. Start: Tue, 25th Jul 2017 15:00 Update: Wed, 26th Jul 2017 18:30 Edited: Wed, 26th Jul 2017 13:31 Status: Partial Maintenance: None



#4914 Zen Web Sites - Planned Maintenance - Secure File Sending (New)

Wed, 26 Jul 2017 11:38:58 +0100

Our engineers will be performing maintenance on our secure file sending facility. The secure file sending facility will be unavailable during the maintenance period while updates are applied. This will prevent you from downloading files we have sent to you. Zen regret any inconvenience this may cause. Start: Sun, 13th Aug 2017 21:00 End: Sun, 13th Aug 2017 23:55 Edited: Wed, 26th Jul 2017 11:38 Status: Outage Maintenance: Planned



#4910 Leased Line - Emergency Capacity Maintenance Manchester Area (Close)

Tue, 25 Jul 2017 22:00:56 +0100

All work was completed successfully. Start: Mon, 17th Jul 2017 23:00 End: Tue, 18th Jul 2017 00:55 Clear: Mon, 17th Jul 2017 22:00 Edited: Tue, 25th Jul 2017 22:00 Status: Up Maintenance: None



#4912 Broadband (xDSL) - Network Outage (Close)

Tue, 25 Jul 2017 15:44:35 +0100

All service have now recovered. The outage was caused by a core network device failure. Start: Tue, 25th Jul 2017 14:00 Clear: Tue, 25th Jul 2017 15:40 Edited: Tue, 25th Jul 2017 15:44 Status: Up Maintenance: None



#4913 Leased Line - Network Outage (Close)

Tue, 25 Jul 2017 15:44:04 +0100

All service have now recovered. The outage was caused by a core network device failure. Start: Tue, 25th Jul 2017 14:00 Clear: Tue, 25th Jul 2017 15:40 Edited: Tue, 25th Jul 2017 15:44 Status: Up Maintenance: None



#4913 Leased Line - Network Outage (New)

Tue, 25 Jul 2017 14:35:47 +0100

Zen engineers are investigating the cause of a network outage at around 14:00 today. Most services have now recovered. Further updates will be posted as they become available. Start: Tue, 25th Jul 2017 14:00 Update: Tue, 25th Jul 2017 15:30 Edited: Tue, 25th Jul 2017 14:35 Status: Partial Maintenance: None



#4912 Broadband (xDSL) - Network Outage (New)

Tue, 25 Jul 2017 14:34:52 +0100

Zen engineers are investigating the cause of a network outage at around 14:00 today. Most services have now recovered. Further updates will be posted as they become available. Start: Tue, 25th Jul 2017 14:00 Update: Tue, 25th Jul 2017 15:30 Edited: Tue, 25th Jul 2017 14:34 Status: Partial Maintenance: None



#4911 Shared Linux Hosting - cPanel rscp01 Server Reboot (Close)

Wed, 19 Jul 2017 14:46:12 +0100

Maintenance completed Start: Wed, 19th Jul 2017 13:05 Clear: Wed, 19th Jul 2017 14:45 Edited: Wed, 19th Jul 2017 14:46 Status: Up Maintenance: None



#4911 Shared Linux Hosting - cPanel rscp01 Server Reboot (New)

Wed, 19 Jul 2017 13:07:18 +0100

Our engineers are currently working on rscp01.hosting.zen.net.uk. As part of this work a server reboot will be taking place and customers may notice a brief outage while this happens. Zen regret any inconvenience this may cause. Start: Wed, 19th Jul 2017 13:05 Update: Wed, 19th Jul 2017 14:00 Edited: Wed, 19th Jul 2017 13:07 Status: Outage Maintenance: Emergency



#4904 Shared Linux Hosting - cPanel Planned Maintenance (Close)

Tue, 18 Jul 2017 00:40:39 +0100

Maintenance successfully completed. Confirmed functioning service. Incident Closed. Start: Wed, 28th Jun 2017 23:00 End: Mon, 17th Jul 2017 21:28 Clear: Tue, 18th Jul 2017 00:40 Edited: Tue, 18th Jul 2017 00:40 Status: Up Maintenance: None



#4910 Leased Line - Emergency Capacity Maintenance Manchester Area (New)

Mon, 17 Jul 2017 19:06:35 +0100

We need to complete works to resolve some capacity issues for our Manchester connections. Circuits that connect to Blackfriars and Manchester Central Exchanges may notice some brief down time whilst the work is ongoing. Zen regret any inconvenience this may cause. Start: Mon, 17th Jul 2017 23:00 End: Tue, 18th Jul 2017 00:55 Edited: Mon, 17th Jul 2017 19:06 Status: Partial Maintenance: Emergency